### v1

# Preparing for Super-Reactivity: Early Fault-Detection in the Development of Exceedingly Complex Reactive Systems

Abstract
We introduce the term Super-Reactive Systems to refer to reactive systems whose construction and behavior are complex, constantly changing and evolving, and heavily interwoven with other systems and the physical world. Finding hidden faults in such systems early in planning and development is critical for human safety, the environment, society and the economy. However, the complexity of the system and its interactions and the absence of adequate technical details pose a great obstacle. We propose an architecture for models and tools to overcome such barriers and enable simulation, systematic analysis, and fault detection and handling, early in the development of super-reactive systems. The approach is facilitated by the inference and abstraction capabilities and the power and knowledge afforded by large language models and associated AI tools. It is based on: (i) deferred, just-in-time interpretation of model elements that are stored in natural language form, and (ii) early capture of tacit interdependencies among seemingly orthogonal requirements.
Index Terms:
Super-Reactive Systems, System Engineering, Software Engineering, LLM, Models, Specifications, Interaction, Simulation, Verification, Autonomous systems, Systems of Systems.
I Introduction
Since the 1985 identification of the category of reactive systems[1], a plethora of methods, languages and tools have been introduced to support the development of such systems. Today, complex reactive systems are penetrating almost every aspect of life, including communications, commerce, finance, healthcare, aviation, land transportation, manufacturing, entertainment, and more. The complexity of new systems is compounded by the fact that they are interwoven with other systems and with the physical world, and are constantly changing and evolving. We term this emerging kind of system super-reactive (SR). While system and software engineering (SySE) is benefitting from new developments in generative AI and large language models (LLMs), the challenge of building safe and reliable SR systems remains open. Despite applying the best tools and methodologies, any given system is likely to conceal undesired and very often even unsafe behaviors and impending failures, with the risk of adverse effects on human life, the environment, society and the economy. Thus, while early discovery and handling of such faults is required, it remains a tantalizing challenge. In this paper, we propose tackling this issue relying on the following principles: (1) Allow model elements that are expressed in natural language (NL), benefitting from the expressive power of natural language, its sensitivity to delicate context variations and its ability to navigate multiple levels of abstraction, and carrying out just-in-time (JIT), deferred, interpretation of such NL elements. (2) Discover and document otherwise-tacit interdependencies among separately specified, seemingly orthogonal requirements.
A key enabler for our approach is the availability of large language models and other AI tools, the power and breadth of which is ever-growing. We exploit these in novel ways, in order to render possible principles (1) and (2) above.
Figure 1: Solution architecture blueprint. See explanation in text.
II The Problem
As the complexity and pervasiveness of reactive systems and systems of systems keeps growing, so do the risks associated with hidden faults: potential failure points, malfunction, undesired emergent behaviors and absence of desired ones. Recent well-known cases involving actual system code, maintenance procedures, interfaces with other systems and with humans, etc., include: the Crowdstrike server failures in July 2024, the 2023 accident in which a robotic taxi hit a pedestrian in San Francisco, the failure of the USA FAA notification system in 2023, and the crashes of 737 MAX airplanes. Similar kinds of problems obviously occur frequently without gaining broad attention. Furthermore, beyond such direct effects, issues with an SR system may inflict excessive rigidity and regulation on the behavior of humans and of other systems, in an effort to accommodate the system’s limitations. Compare for example the assignment of cars to driving sides and lanes on roads; one would like to avoid having such restrictions enforced in sidewalks and hallways shared by humans and robots as an emergency response to an unanticipated problematic reality.
Published discussion of such problems call for early assessment and preemptive technical, economic, and regulatory activities. Early in the days of model-based system engineering (MBSE), France and Rumpe wrote: It is our view that software engineering is inherently a modeling activity, and that the complexity of software will overwhelm our ability to effectively maintain mental models of a system. [2]. Over the years there was great progress in the ability to build executable models. Examples include UML, SysML, Rhapsody, STATEMATE, fUML, xUML, Ptolemy, MATLAB with Simulink and Stateflow, SCADE, UPPAAL, BPMN, Arcadia, Cameo and others; see also list of SySE tools in  [3, pp.76,79,208]). In parallel, there were significant advances in applying formal methods to such models [4, 5, 6, 7, 8, 9, 10, 11, 12]. And, many organizations developed elaborate ad-hoc models to help study the systems early in, and throughout, development [13, 14, 15].
However, despite such advances, ensuring the safety and correctness of complex systems is still an unsolved problem. For example, in a 2024 workshop on safety of autonomous transportation (summarized in [16]) many open SySE issues and challenges were discussed, including: (i) incorporating general and domain knowledge in testing and verification; (ii) ensuring that ML training data covers rare but critical scenarios; (iii) exhaustively covering all possible interactions; and, (iv) enhancing usability of formal methods. Similar conclusions about gaps in present methods for early issue identification appear in [17, 18, 19, 20]. Given that uncovering hidden faults in well-specified and even fully developed systems is still an open problem, it is evident that preemptive fault discovery in super-reactive systems (e.g., complex systems of systems interwoven with their environment) at early development stages poses a major challenge. The added difficulty stems in part from the informal and imprecise nature of requirements in early development stages, the limited scalability of the tools, and reliance on engineers to infer at development time undocumented relations among separately specified requirements that use different abstractions and a variety of terminologies, the dependence of verification on translation to state machines and Petri nets, and the absence of executable and analyzable semantics of certain specification artifacts.
Current AI-based solutions assist in various activities of development, including code generation to debugging, with a prominent example being the GitHub Copilot; see, also, e.g., [21] and references therein. However, applying such tools in the context of early specification is mostly limited to automated modeling, which we discuss in Section III-C.
The roadmap presented here addresses these issues by enabling rigorous execution and analysis subject to domain expertise and world knowledge, and doing so at higher levels of abstraction. The approach relies on the ability of AI-based tools to mimic humans’ flexible navigation of complex abstraction relations using natural language, and will extend the present use of abstraction as in object-oriented inheritance relations, and in counter-example guided abstraction refinement (CEGAR) in formal verification [22, 23].
III The roadmap
Below we list the elements of an approach and an architecture for modeling SR systems, including a set of intelligent tools for simulation and analysis, which, together, can enable the much desired early preemptive discovery of hidden faults, while coping with the existing challenges and technology gaps. In bounding the scope of the proposed solution, we exclude the use of AI, ML and LLMs in runtime decision making, monitoring, DevOps, or the formal verification of final code . And, while some reasoning functions of the proposed solutions may be similar to those carried out by expert human engineers and domain professionals, we focus on enabling presently impractical or impossible analyses, and much less on automation of manual tasks.
III-A The Intelligent Development Aide
The Intelligent Development Aide (IDA) is a shared layer of services that offers the following to the overall solution: (i) intelligence, including learning, inference, and generative abilities; (ii) NL-based interaction capabilities; and, (iii) general world knowledge and certain domain-specific expertise. The IDA will rely on present and future technologies and tools that come under the umbrella of AI, Generative AI, Machine Learning, Deep Learning, Large Language Models (LLMs), etc. It will be constructed, among other things, by fine-tuning, enhancing and extending AI-based tools, relying on techniques like those of  [24, 25, 26, 27, 28] and future emerging ones. With inputs coming from diverse systems, with textual and visual depictions of execution scenarios and logs of normal and faulty interactions, the IDA will be trained to recognize unique software and system engineering issues and new delicate kinds of interdependencies.
III-B The i-model
We introduce a new kind of model, termed i-model, which offers fresh perspectives on some common modeling maxims:
First, while precision is commonly needed to ensure correct system implementation, i-models will take advantage of what may appear quite the opposite. They will retain within model entities the expressive power of NL, which includes sensitivity to context, flexible abstraction, generalization, associations, etc. Simulation and analysis tools will then rely on deferred -– just-in-time (JIT) -– interpretation to endow NL and NL-like behavioral specifications with concrete meaning, aligned with the intended context, and abstraction level.
Second, while logical flow and organization are essential to engineering, conceptual abstractions may not always lend themselves to being so depicted. For example, consider the difficulty in modeling a complex network of multiple class inheritances, combined with natural language ambiguity, where, for instance, the word stop could mean a condition of no motion at all, or a process of slowing down to reach that condition, or the action of beginning to press the brake in order to begin this process, etc. In contrast, i-models will accommodate coexistence of multiple, diverse, non-hierarchical, overlapping and dynamic abstraction lattices.
And third, modularity, encapsulation, and logical decomposition are central principles in software engineering, and in engineering in general. However, separately specified requirements often have tacit, unstated dependencies, which show up as exceptions, priorities, alternatives, complementary or concurrent actions, mutually exclusive conditions, etc. It was up to the engineers to infer these implicit relations, and to reflect their understanding in the implementation. In our automated construction of i-models from a wide range of specifications, special focus will be put on discovering such unstated relationships and capturing them in the model, despite the entanglement that they may imply.
The i-models will store diverse information, including requirements, goals, behaviors, scenarios, properties, emergent effects, etc., as well as groupings, abstractions, and relations of such entities. It will also contain meta information about potential changes due to evolution of the system and its environment, allowing further analysis of potential future trajectories.
Finally, the i-model will support unmodeling; i.e., capturing entities and assumptions that should be ignored during execution and analysis, as well as operational environments in which the SR system is not expected to operate. Unmodeling will complement capabilities in existing modeling techniques to specify the exact intended operational design domain (ODD), directing IDA-based tools where to apply and where not to apply their vast and important knowledge [29].
The management of the immense knowledge stored in the i-model will be divided among three realms: (i) the entities themselves, including structured data and unstructured documents, much of which can be in NL; (ii) the rich relationship between entities, which will be represented in the i-model database; and, (iii) the general and application-specific knowledge captured in the IDA components, both in advance, and following the training associated with the building and analysis of the i-model.
III-C The i-model builder
Inputs to i-model building will include: requirements documents, specifications of reusable components [30], manual risk analyses [31, 32], whole models in various modeling languages, program code, documentation, example run logs of early prototypes, test cases, etc. Additional information, corrections and guidance provided interactively by engineers during model building will also be retained. Furthermore, the i-model builder can initiate queries, soliciting the engineers for missing information or confirmation of inferences.
Beyond the now increasingly common translation of NL specifications into basic object models and computer programs, a unique feature of the i-model builder will be the automated, and optionally interactive, discovery and recording of undocumented tacit interdependencies among separately specified entities. For example, consider separately specified rules for an autonomous vehicle (AV) that may cause the AV to accelerate, as when entering a highway, or when instructed to follow another vehicle, or when returning to normal speed after a temporary slow down; consider a second set of rules specifying maximum legal speed and maximum recommended speed under certain conditions. The fact that the rules in the second set constrain or may be in conflict with rules in the first set, will be captured at model-building time. When the effect is clear, i.e., that one rule takes priority over the other, this explicit specification will be generated automatically up front (where today it may be left as an implementation detail); when the relation is in question, (say, what to do if the leader of a convoy exceeds the legal speed limit), the i-model builder will consult the engineers.
The builder will provide succinct summaries of the input information, and elaborate on its inferences applying logic and domain knowledge. The added information will also be stored in the i-model. For example, in our experiments with an LLM in creating a model from a description of a traffic scenario, the LLM added pedestrian objects, which were absent from the original requirements.
The i-model will use a rich modern database to store the structured and unstructured information and the associated connections and relationships. As this structure will likely be too complex for humans to navigate and maintain directly, it will be supported by a multi-view presentation (see, e.g., slices in [33]), projecting requested information as stand alone succinct text descriptions and dynamically created diagrams (e.g., Statecharts). Since many entities will already be in NL, the projection itself will be intuitive, but may still require other components for confirmed interpretation. Once confirmed, such projections can also serve as concise memory, to be fed back to the i-model.
To illustrate related capabilities of AI tools we conducted several experiments. For example, we described to an LLM requirements and properties for the edges and nodes of graphs. The LLM was able to produce the full list of nodes and edges; its few mistakes highlighted omissions and ambiguity in the specification. For example, we prompted the LLM as follows: “… Consider the set S containing the entities A, B, C, D. Consider all subsets A1-A15 of S containing at least one element. … Consider the following graph. From each subset A_i there is an undirected edge to each subset that has fewer entities than A_i (they do not have to be A_i’s subsets). [Furthermore], all the subsets with exactly 2 elements are connected to each other with directed edges in a cycle - i.e., each one has exactly one entering arrow and one exiting arrow. The order does not matter…”.
The LLM’s response consisted of a complete list of nodes, followed by the list of edges, including, for example, {A,B}: Directed edge to {A,C}; {A,C}: Directed edge to {A,D}”. We then asked it to describe a higraph [34] having four basic nodes, blobs around certain sets of nodes, and certain edges between blobs. To make the problem real, we cast the blobs as assignments of persons to projects, and the edges as dependencies among projects. The LLM was able to create a bulleted list of all the blobs and their contents, and connecting edges.
We expect the construction of the i-model builder to extend the growing amount of research on using GAI and LLMs for creating models from NL inputs for; see, e.g.  [35, 28, 36, 37, 38, 39]
III-D The i-model execution and simulation engine
This component carries out dynamic, deferred, just-in-time (JIT) interpretation of the definite and the NL information in the i-model, and outputs descriptions of possible runs and the unfolding of “what if” scenarios at desired abstraction levels, in the forms of narratives, event logs, and even synthetic videos.
For illustration, consider the following example, which we ran on OpenAI ChatGPT. Prompt (slightly edited): “Consider a 3 lane street. 30 equal-size passenger cars are filling the street, standing in front of a traffic light. No cars arrive from the back of the queue throughout this scenario. The third car stalls and cannot move. No one stops to help. Describe the behaviors and positions of the 30 cars after the light turns green. Assume the light is green for a very long time.”.
Here are a few excerpts from the LLM’s response: “When the traffic light turns green: …
1. Cars in the Middle Lane (Lane 2): (a) The first two cars in the middle lane (Cars 1 and 2) will start moving forward as the light turns green. (b) The third car in the middle lane (Car 3) will not move… (c) The cars behind the third car … will initially wait, expecting the third car to move. After realizing that it is not moving, they will either: (i) Start honking, …; (ii) Some drivers may try to change lanes … (iii) Some drivers may remain stuck behind the third car if changing lanes is not possible…
2. Cars in the Left Lane… and Right Lane…: (a) The cars in these lanes … will begin moving forward … (b) The first few cars … positioned parallel to the stationary third car, will drive past it without much issue. (c) The cars behind them will continue moving forward, potentially picking up speed…”
When we divided the road into small virtual cells, the LLM was able to describe car movements through the numbered cells, making the textual scenario more concrete and mathematically analyzable.
The LLM responses also contained errors; for example, the LLM suggested the possibility that cars blocked behind the stalled car will not be able to change lanes at all, ignoring the specified facts that the traffic light stays green for a long time, and that no new cars arrive during the scenario in question; however, we believe that with proper training, responses can be made highly reliable. And in any case, the answers can be checked and validated interactively, and domain-specific knowledge can be added. Note how the LLM incorporated into the responses its domain knowledge; for example, the adding the possibilities of changing lanes and honking, which were not included in the original specification.
The execution engine will, of course, benefit from state-of-the-art execution and simulation techniques, like those in SysML, Statecharts, and Scenario-based programming[40, 41], as well as in direct execution of NL specifications as in [27] and references therein.
Borrowing from techniques for test-case generation [42], execution and simulation engine will also generate and process batches of diverse, yet relevant, “what if” scenarios, and store their execution results for further processing.
III-E The i-model analysis engine
The i-model analyzer will carry out the equivalent of formal model-checking, searching — proactively — for execution trajectories that lead to fault states. Treating the model as an NL-enriched graph, it will traverse its paths, interpreting entities and relationships subject to general and domain knowledge, including causalities, interdependencies and risks, while abiding by unmodeling — knowing what to exclude. In addition, the analyzer will offer query capabilities, e.g., for investigating complex scenarios, or the many connections of a given entity. It will also interface with classical model checkers and satisfiability modulo theory (SMT) constraint solvers for inspecting well-structured subsets and projections of the i-model, and for presenting the answers back in NL.
In our experiments, an LLM was able to discover paths and absence thereof in graphs like those built in the example of Section III-C. Other analytic LLM capabilities that can support i-model analysis are described in papers like [37]. These include computing when two independent periodic events may occur simultaneously, explaining behavior, articulating system properties, etc. Furthermore, it is expected that LLM general and domain-specific analytical capabilities will be extended and deepened, and are likely to be intertwined with ongoing research in software and system engineering. Developments along the present roadmap can incorporate such enhancements, and target them specifically at early fault-detection.
III-F Emergent Effect Detector
This component accepts outputs of system simulations, looking for expected and unexpected patterns and emergent effects, both structural/spatial and behavioral/temporal. Such effects may be previously specified as desired, undesired, or perhaps acceptable, or they may require assessment. The tool will rely on the immense body of work in recognizing patterns, emergent effects, anomalies, etc., in sequential data, like discrete event logs or continuous signals, and in spatial and structural information, like images and videos. See, e.g.,  [43, 44, 45, 46]. The results will be presented formally and in NL for manual and automated analysis.
III-G Repair Advisor
The i-model’s sheer size may interfere with its maintenance, calling for a repair advisor that accepts a description of an issue and proposes changes to the system or to its technical and physical environment. A key distinction from common program repair [47] is the primary focus on pinpointing the model components that should be changed and on describing the ensuing impact on system behavior, while the technical details of the actual change are secondary.
IV Conclusion
We are now in the process of initiating a research project following the above roadmap. Development of models and tools that enable simulation and analysis of highly complex systems based only on early specifications can dramatically enhance our ability to develop reliable, safe, and productive SR systems. A combination of the recent advances in AI and a fresh perspective on what may and may not qualify as a model entity or be acceptable as a simulation result, may enable the achievement of this tantalizing goal.
Acknowledgments
This research was funded in part by an NSFC-ISF grant to DH, issued jointly by the National Natural Science Foundation of China (NSFC) and the Israel Science Foundation (ISF grant 3698/21). Additional support was provided by a research grant to DH from Louis J. Lavigne and Nancy Rothman, the Carter Chapman Shreve Family Foundation, Dr. and Mrs. Donald Rivin, and the Estate of Smigel Trust.